3y(dy/ dx)=-3x ; y(2)=-π unique or not?
1
Expert's answer
2019-09-11T13:39:55-0400

We have the equation "\frac{dy}{dx}=-\frac{x}{y}" with initial condition "y(2)=-\pi". One of solutions for this equation is "x^2+y^2=\pi^2+4" . It can be transformes into two functions: "y(x)=\sqrt{\pi^2+4-x^2}" and "y(x)=-\sqrt{\pi^2+4-x^2}" .Function "y(x)=\sqrt{\pi^2+4-x^2}" does not satisfy the initial condition "y(2)=-\pi" . The second function,

"y(x)=-\sqrt{\pi^2+4-x^2}" , defined on "(-\sqrt{\pi^2+4},\sqrt{\pi^2+4})", satisfies the initial condition.

From https://www.science.unitn.it/~bagagiol/noteODE.pdf, theorem 2.10 we have that for all "a\in (0,\pi)" the equation has a unique local solution in "(-\sqrt{\pi^2+4},\sqrt{\pi^2+4})\times (-\infty,a)" (which is "y=-\sqrt{\pi^2+4-x^2}" ).

Hence the equation has a unique local solution on "(-\sqrt{\pi^2+4},\sqrt{\pi^2+4})\times (-\infty,0)" .Indeed, assume that there is solution "y_1(x)" defined on "I" , where "y_1(u)\neq -\sqrt{\pi^2+4-u^2}" for some "u\in (-\sqrt{\pi^2+4},\sqrt{\pi^2+4})\cap I" . If"v=\max\{y_1(u), -\sqrt{\pi^2+4-u^2}\}" , we have two different solutions in "\bigl(I\cap (-\sqrt{\pi^2+4},\sqrt{\pi^2+4})\bigr)\times (-\infty,\frac{v}{2})" . Contradiction with uniqueness of "y(x)=-\sqrt{\pi^2+4-x^2}".

Assume that there is solution "y_2(x)", defined on "I_1" such that "I_1\setminus [-\sqrt{\pi^2+4},\sqrt{\pi^2+4}]\neq\varnothing" . If there is "c\in I_1\cap (-\infty,-\sqrt{\pi^2+4})" , then "[c,2]\subset I_1" and "y_2(x)=-\sqrt{\pi^2+4-x^2}" on "[-\sqrt{\pi^2+4},2]" . So "y_2(-\sqrt{\pi^2+4})=0" and "\bigl|y_2'(-\sqrt{\pi^2+4})\bigr|=+\infty".

Let "x\in (c,-\sqrt{\pi^2+4})". Since "y_2'(x)=-\frac{x}{y_2(x)}" and "x<0" , we have "sgn (y_2'(x))=sgn(y_2(x))" .

If "y_2'(-\sqrt{\pi^2+4}-0)=-\infty" , then "y'(x)<0" in "(d,-\sqrt{\pi^2+4})" for some "d\in (c,-\sqrt{\pi^2+4})" . But since "y_2" is decreasing function in "(d,-\sqrt{\pi^2+4})" and "y_2(-\sqrt{\pi^2+4})=0" , we have "y_2>0" in "(d,-\sqrt{\pi^2+4})" . Contradiction with statement that "sgn (y_2')=sgn(y_2)" in "(d,-\sqrt{\pi^2+4})" . Similarly we obtain a contradiction while assuming that "y_2'(-\sqrt{\pi^2+4}-0)=+\infty". So "I_1\cap (-\infty,-\sqrt{\pi^2+4})=\varnothing" .

Likewise we can obtain that "I_1\cap (\sqrt{\pi^2+4},+\infty)=\varnothing" .

Hence every solution of the equation is defined on a subset of "[-\sqrt{\pi^2+4},\sqrt{\pi^2+4}]" . Uniqueness of the local solution of "y(x)=-\sqrt{\pi^2+4-x^2}" in "[-\sqrt{\pi^2+4},\sqrt{\pi^2+4}]\times (-\infty,0]" gives us uniqueness of the solution "y(x)=-\sqrt{\pi^2+4-x^2}" in "\mathbb R^2" .